OpenStudy (anonymous):

What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = –1/3x – 2?

OpenStudy (anonymous):

Look at your slope of that line. The slope is -1/3. In order for your other line to be perpendicular, it needs to be the "negative inverse." That means you have to flip the -1/3 and then multiply it by -1. -1*(-3/1) = 3 So you have your point (12,4) and your slope m=3 Plug these values into y=mx+b to get your b-value (y-intercept) After that, you'll have your m-value and your b-value. Plug those (only those) into y=mx+b and you'll have the eqn of your perpendicular line.

OpenStudy (anonymous):

-4, 6

OpenStudy (anonymous):

Okay, I'll start from here: So you have your point (12,4) and your slope m=3 Plug these values into y=mx+b to get your b-value (y-intercept) 4 = 3*(12) +b b=4- (3)*12=-32 m=4 b=-32 Plug them into y=mx+b and you have the equation of your line. You're not looking for a point. You're finding an equation.

OpenStudy (anonymous):

Do you know how to plug numbers in to formulas?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Read through this whole thing then. You won't be able to do these problems if you don't know how to graph or plug numbers in. http://tutorial.math.lamar.edu/Classes/Alg/Lines.aspx